Circular No. 6698

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cv IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (V = 19.7, R = 19.4) on CCD images taken on July 7 by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cv is located at α = 13^{h}35^{m}20^{s}.77, δ = −34^{o}12′11′′.6 (equinox 2000.0), apparently in a large spiral galaxy near Abell 3565. A nearby star has position end figures 26^{s}.38, 11′46′′.6. SN 1997cv was also present, and comparably bright, on similar frames taken on July 10, but it is not visible on images taken on June 10 (limiting R ∼ 21). Finding charts are available at http://msowww.anu.edu.au/∼reiss/

Abell SNSearch/.

CEPHEUS X-4 P. Roche and L. Green, University of Sussex; and M. Hoenig, Institute

of Astronomy, Cambridge, report: “We propose the identification of the transient 66-s x-ray pulsar Cep X-4 = GS 2138+56 (IAUC 2493, 2512, 4575, 4577) with a V = 14.2 Be star, located at α = 21^{h}39^{m}30^{s}.6, δ = +56^{o}59′12′′.9 (equinox 2000.0), based on positional coincidence with the ROSAT error circle of Schulz, Kahabka and Zinnecker (1995, A.Ap. 295, 413). The object displays Balmer (Hα equivalent width = 5.3 nm) and Fe II emission on a B0 or earlier spectrum, indicating a strong Be phase. Given current indications of x-ray activity from RXTE ASM, optical and infrared monitoring are encouraged.”

COMET C/1997 J1 (MUELLER) Total visual magnitude estimates: May 30.97 UT, 13.2 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); June 3.90, 12.9 (M. Lehky, Hradec Kralove, Czech Republic, 0.42-m reflector); 7.92, 12.7 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 9.91, 11.9 (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector); 29.50, 13.6 (A. Nakamura, Kuma, Ehime, Japan, 0.6-m reflector + CCD).

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: June 10.93 UT, 13.3 (Lehky), 23.93,

11.9 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 28.90, 12.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); July 3.95, 12.2 (C. Vitorino, Cabo da Roca, Portugal, 0.25-m reflector); 6.11, 12.4 (J. E. Bortle, Stormville, NY, 0.41-m reflector).

1997 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Timothy B. Spahr